### v1

@@ Abstract
Our interest lies in developing some efficient methods for minimizing the sum of two geodesically convex functions on Hadamard manifolds, with the aim to enhance the convergence of the Douglas-Rachford algorithm in Hadamard manifolds. Specifically, we propose two types of algorithms: inertial and non-inertial algorithms. The convergence analysis of both algorithms is provided under suitable assumptions on algorithmic parameters and the geodesic convexity of the objective functions. This convergence analysis is based on fixed-point theory for nonexpansive operators. We also study the convergence rates of these two methods.
@@ 1 Introduction
The Douglas-Rachford algorithm is among the most celebrated splitting algorithms in convex analysis and operator theory [1, 2, 3, 4, 5, 6], originally introduced by Douglas and Rachford [7] for the numerical solution of partial differential equations. It was later generalized by Lions and Mercier [8] and independently by Passty [9] to solve the more general problem of finding a zero of the sum of two maximally monotone operators in Hilbert spaces. Eckstein and Bertsekas [10] provided further insights into its convergence under relaxation schemes and inexact computations.
Beyond operator inclusions, the algorithm also finds wide application in convex optimization, particularly for problems involving the sum of two proper, lower semicontinuous, and convex functions; see [1]. Given a Hilbert space $X$ and two proper, convex, lower semicontinuous functions $\varphi,\psi:X\to(-\infty,+\infty]$, the Douglas–Rachford algorithm aims to solve the following convex minimization problem:
$\min_{x\in X}\varphi(x)+\psi(x).$ (1)
The strength of the Douglas-Rachford algorithm lies in its splitting structure, which allows one to compute the proximal mappings of $\varphi$ and $\psi$ separately. This structure makes the algorithm especially useful in large-scale and structured optimization problems [1].
Recently, these algorithms were successfully applied in image processing mainly for two reasons: the functional to minimize allows simple proximal mappings within the method and it turned out that the algorithms are highly parallelizable; see [1]. Therefore, Douglas-Rachford methods became some of the most popular variational methods for image processing. Continued interest in Douglas-Rachford iteration methods is also due to its excellent performance on various nonconvex problems; see [11, 12].
Motivated by applications where data naturally reside on nonlinear spaces such as manifolds, common in computer vision, machine learning, and information geometry, there has been growing interest in extending convex optimization methods to manifold settings. Hadamard manifolds, in particular, offer a natural framework for such generalizations due to their geometric properties, including unique geodesics between points [13]. In the past two decades, several results from linear spaces have been extended to manifold settings because of their applications in many areas of science, engineering, and management (see [14, 15, 16, 17]).
@@
$\min_{x\in\mathcal{M}}\varphi(x)+\psi(x).$ (2)
To compute the solution to problem (2), Bergmann et al. [4] proposed the following Douglas-Rachford algorithm on Hadamard manifold:
@@
where $\{\alpha_{n}\}$ is a sequence in $[0,1]$ with $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$ and $\lambda>0$. Bergmann et al. [4] showed that if $R_{\lambda\phi}R_{\lambda\psi}$ is nonexpansive, then the sequence generated by method (3) converges to an element $v$ of $\operatorname*{Fix}(R_{\lambda\phi}R_{\lambda\psi})$ such that $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2).
It is worth noting that by setting $T={R}_{\lambda\phi}{R}_{\lambda\psi}$, the Douglas-Rachford algorithm (3) can be seen as a Krasnoselski-Mann iteration given by
@@
where $\{\alpha_{n}\}$ is a sequence in $[0,1]$ with $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$ and $\lambda>0$. This observation connects optimization and fixed point theory and hence problem (1) can now be viewed as a fixed point problem.
This naturally leads to the framework of fixed point theory. Various problems across diverse domains, such as image reconstruction [18], signal processing [18], variational inequality problems [19], and convex feasibility problems [20] can be formulated as fixed point problems. Let $C$ be a nonempty, closed and convex subset of a real Hilbert space $X$ and $T:C\to C$ be a nonlinear operator. Then, the fixed point problem can be stated as:
@@
To solve such problems, numerous iterative techniques have been developed. One of the most widely used is the Mann iterative method, introduced by Mann [21] in 1953. The Mann iteration method has its limitations, as it may not effectively approximate the fixed points of pseudocontractive mappings (see [22]). To tackle this challenge, Ishikawa [23] proposed an iterative approach, extensively investigated by multiple authors across different spaces to compute the fixed point of pseudocontractive and nonexpansive mappings (see [24, 25, 26]).
Building on this, Agarwal et al. [27] introduced the S-iteration method as follows:
@@
where $\{\alpha_{n}\}$ and $\{\beta_{n}\}$ are sequences in $(0,1)$ with $\sum_{n=1}^{\infty}\alpha_{n}\beta_{n}(1-\beta_{n})=\infty$. In 2011, Sahu [28] proposed a modified version of the S-iteration method (5), called the normal S-iteration method, also known as the hybrid Picard-Mann iteration method, which is defined as follows:
@@
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$. Continuing in this direction, Sahu [29] observed that applying the operator multiple times in each iteration can significantly enhance the convergence. This led to the $p$-accelerated normal S-iteration method [29], defined as follows:
@@
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ and $p$ is the power of acceleration. For $p=1$, the $p$-accelerated normal S-iteration method reduces to the normal S-iteration method defined by (6). It is observed in [29] that the $p$-accelerated normal S-iteration method has better convergence performance than the Mann iterative method and normal S-iterative method for the fixed point of nonexpansive operators with convex domain in infinite dimensional Hilbert space. In fact in most cases $p$-accelerated normal S-iteration method works better than the inertial Mann iteration method due to the parameter $p$.
Inertial-type extrapolation techniques play a critical role in accelerating iterative methods. The foundation of such methods can be traced back to Polyak [30], who introduced the heavy ball method. Alvarez and Attouch [31] extended the concept to improve the performance of the proximal point algorithm, proposing the inertial proximal point algorithm for solving inclusion problems.
In the existing literature, the utilization of inertial extrapolation-based methods has been thoroughly explored and implemented in various contexts (see [32, 33, 34, 35, 36] and references therein). Recently, researchers have developed several iterative algorithms by incorporating inertial extrapolation techniques, such as the fast shrinkage thresholding algorithm [37], inertial forward-backward algorithm [38] and inertial Douglas-Rachford splitting algorithm [2].
Building upon the development of inertial and accelerated iterative methods, Mainge [39] proposed a novel approach in 2008 by combining inertial extrapolation with the Mann method, aiming to compute the fixed point of nonexpansive mappings on a real Hilbert space. The inertial Mann algorithm is defined as follows:
@@
where $\{\theta_{n}\}$ and $\{\alpha_{n}\}$ are sequences in $[0,1)$. Mainge [39] studied the weak convergence of inertial Mann algorithm (8) under the conditions that $\theta_{n}\in[0,\theta]$ for all $n\in\mathbb{N}$, where $\theta\in[0,1)$, $\sum_{n=1}^{\infty}\theta_{n}\|x_{n}-x_{n-1}\|^{2}<\infty$, and $\inf_{n\in\mathbb{N}}\alpha_{n}>0$ and $\sup_{n\in\mathbb{N}}\alpha_{n}<1$.
The second condition that $\sum_{n=1}^{\infty}\theta_{n}\|x_{n}-x_{n-1}\|^{2}<\infty$, taken by Mainge [39] is often difficult to verify in practice. Boţ et al. [2] proposed a modification in algorithm (8) by replacing this condition with an alternative criterion that the sequence $\{\theta_{n}\}$ is nondecreasing in $0,\theta]$ with $\theta_{1}=0$ and $\theta\in[0,1)$ such that
@@
Recently, Dixit et al. [40] proposed the inertial normal S-iteration method for the solution of problem (4) for a nonexpansive operator, which in practice converges faster than the Mann iteration method, inertial Mann iteration method and normal S-iteration method.
The verification of condition (9), as considered by [2], is a challenging task. To address this issue, Sahu [29] proposed a new convergence criterion for the inertial Mann iteration. This criterion assumes that the sequences $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ satisfy the following conditions: there exist constants $0<a\leq\alpha_{n}\leq b<1$ for all $n\in\mathbb{N}$, the sequence $\{\theta_{n}\}\subset[0,1)$ is nondecreasing, and there exists $\theta\in(0,1)$ such that $0\leq\theta_{n}\leq\theta<\frac{\epsilon}{1+\epsilon+\max\{1,\epsilon\}}$ for all $n\in\mathbb{N}$, where $\epsilon=\frac{1-b}{b}$. These conditions are easier to verify in practice.
Motivated by the work of Sahu [29] and Bergmann et al. [4], the main purpose of this paper is to introduce an inertial Douglas-Rachford method and $p$-accelerated normal S-Douglas-Rachford method for solving the problem (2) and study its convergence analysis. We give application of these algorithms to minimizing functionals containing multiple summands by introducing a parallel inertial Douglas-Rachford method and parallel $p$-accelerated Douglas-Rachford method. We also apply our parallel Douglas-Rachford methods to solve the generalized Heron problems and provide numerical experiments.
@@ 2 Preliminaries
Let $C$ be a nonempty subset of a metric space $(X,d)$ and let $T\colon C\to C$ be an operator. Then $T$ is said to be nonexpansive if
$d(Tx,Ty)\leq d(x,y)\text{ for all }x,y\in C.$
We denote the set of fixed points of an operator $T\colon C\to C$ by $\operatorname*{Fix}(T)$, thus,
$\operatorname*{Fix}(T)=\{x\in C\colon Tx=x\}.$
We denote the set $\mathbb{N}\cup\{0\}$ by $\mathbb{N}_{0}$.
@@
$R_{T,\{x_{n}\}}(n)=\min_{1\leq i\leq n}d(x_{i},Tx_{i})\text{ for all }n\in\mathbb{N}.$ (10)
Observe that $R_{T,\{x_{n}\}}(n)\leq d(x_{i},Tx_{i})$ for all $i=1,2,\ldots,n$ and $\{R_{T,\{x_{n}\}}(n)\}$ is decreasing.
We use the notation $o$ means that $s_{n}=o(1/t_{n})$ if and only if $\lim_{n\to\infty}s_{n}t_{n}=0$.
Definition 1 ([41]).
Let $C$ be a nonempty subset of a metric space $(X,d)$ and let $T\colon C\to C$ be an operator with $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ a sequence in $C$. We say that $\{x_{n}\}$ has
@@
    the LEAF point property if $\{x_{n}\}$ has both LE property and AF property.
Lemma 2 ([29]).
@@
$a_{n+1}\leq a_{n}+\theta_{n}(a_{n}-a_{n-1})+b_{n}\text{ for all }n\in\mathbb{N},\sum_{n=1}^{\infty}b_{n}<\infty$
and there exists a real number $\theta$ with $0\leq\theta_{n}\leq\theta<1$ for all $n\in\mathbb{N}$. Then, the following hold:
@@
For the basic definitions and results regarding the geometry of Riemannian manifolds, we refer to [43, 44, 45].
Let $\mathcal{M}$ be an $m$-dimensional differentiable manifold and let $x\in\mathcal{M}$. The set of all tangent vectors at the point $x$, denoted by $\mathcal{T}_{x}\mathcal{M}$, is called the tangent space of $\mathcal{M}$ at $x$. It forms a real vector space of dimension $m$. The tangent bundle of $\mathcal{M}$ is defined as the disjoint union $\mathcal{T}\mathcal{M}=\bigcup_{x\in\mathcal{M}}\mathcal{T}_{x}\mathcal{M}$, which is a smooth manifold of dimension $2m$.
We suppose that $\mathcal{M}$ is equipped with a Riemannian metric, i.e., a smoothly varying family of inner products on the tangent spaces $\mathcal{T}_{x}\mathcal{M}$. The metric at $x\in\mathcal{M}$ is denoted by $\langle\cdot,\cdot\rangle_{x}:\mathcal{T}_{x}\mathcal{M}\times\mathcal{T}_{x}\mathcal{M}\to\mathbb{R}$. A differentiable manifold $\mathcal{M}$ with a Riemannian metric $\langle\cdot,\cdot\rangle$ is said to be a Riemannian manifold. The induced norm on $\mathcal{T}_{x}\mathcal{M}$ is denoted by $\|\cdot\|_{x}$.
Let $x,y\in\mathcal{M}$ and let $\gamma\colon[a,b]\to\mathcal{M}$ be a piecewise smooth curve joining $x$ to $y$. Then the length of the curve $\gamma$ is defined by $L(\gamma)\coloneq\int_{a}^{b}\|\gamma^{\prime}(t)\|dt$. The minimal length of all such curves joining $x$ to $y$ is called the Riemannian distance, denoted by $d(x,y)$, i.e., $d(x,y)\coloneq\inf\{L(\gamma)\colon\gamma\in C_{xy}\}$, where $C_{xy}$ denotes the set of all piecewise smooth curves $\gamma\colon[0,1]\to\mathcal{M}$ such that $\gamma(0)=x$ and $\gamma(1)=y$.
A single-valued vector field on a differentiable manifold $\mathcal{M}$ is a smooth mapping $V\colon\mathcal{M}\to T\mathcal{M}$ such that, for each $x\in\mathcal{M}$, $V(x)\in\mathcal{T}_{x}\mathcal{M}$. Let $\nabla$ be the Levi-Civita connection associated with the Riemannian manifold $\mathcal{M}$. A vector field $V$ along $\gamma$ is said to be parallel if $\nabla_{\gamma^{\prime}}V=0$. If $\gamma^{\prime}$ is parallel along $\gamma$, i.e., $\nabla_{\gamma^{\prime}}\gamma^{\prime}=0$, then $\gamma$ is said to be geodesic. In this case, $\|\gamma^{\prime}\|$ is a constant. Moreover, if $\|\gamma^{\prime}\|=1$, then $\gamma$ is called a unit-speed geodesic. A geodesic joining $x$ to $y$ in the Riemannian manifold $\mathcal{M}$ is said to be a minimal geodesic if its length is equal to $d(x,y)$. We denote geodesic joining $x$ and $y$ by $\gamma(x,y;\cdot)$, i.e., $\gamma(x,y;\cdot)\colon[a,b]\to\mathcal{M}$ is such that, for $a,b\in\mathbb{R}$, we have $\gamma(x,y;a)=x$ and $\gamma(x,y;b)=y$.
@@
The parallel transport on the tangent bundle $\mathcal{T}\mathcal{M}$ along the geodesic $\gamma$ with respect to Riemannian connection $\nabla$ is defined as $P_{\gamma,\gamma(b),\gamma(a)}\colon\mathcal{T}_{\gamma(a)}\mathcal{M}\to\mathcal{T}_{\gamma(b)}\mathcal{M}$ such that $P_{\gamma,\gamma(b),\gamma(a)}(v)=V(\gamma(x,y;b))\text{ for all }a,b\in\mathbb{R},v\in\mathcal{T}_{x}\mathcal{M},$ where $V$ is the unique vector field satisfying $\nabla_{\gamma^{\prime}(x,y;t)}V=0$ for all $t\in[a,b]$ and $V(\gamma(x,y;a))=v$. A Riemannian manifold $\mathcal{M}$ is said to be complete if for any $x\in\mathcal{M}$ all geodesic emanating from $x$ are defined for all $t\in\mathbb{R}$. By Hopf-Rinow Theorem [44], if $\mathcal{M}$ is complete, then any pair of points in $\mathcal{M}$ can be joined by a minimal geodesic. Moreover, $(\mathcal{M},d)$ is a complete metric space. If $\mathcal{M}$ is a complete Riemannian manifold, then the exponential map $\exp_{x}\colon\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ at $x\in\mathcal{M}$ is defined by $\exp_{x}v=\gamma_{x,v}(1)$ for all $v\in\mathcal{T}_{x}\mathcal{M}$, where $\gamma_{x,v}\colon\mathbb{R}\to\mathcal{M}$ is a unique geodesic starting from $x$ with velocity $v$, i.e., $\gamma_{x,v}(0)=x$ and $\gamma^{\prime}_{x,v}(0)=v$. It is known that $\exp_{x}(tv)=\gamma_{x,v}(t)$ for each real number $t$. For the exponential map, we have the following assertions:
@@
  4. (iv)
    Let $y\in\mathcal{M}$ and $v=\exp_{x}^{-1}y\in\mathcal{T}_{x}\mathcal{M}$. Then $\exp_{x}(tv)=\exp_{x}(t\exp_{x}^{-1}y)=\gamma_{x,\exp_{x}^{-1}y}(t)$ for all $t\in\mathbb{R}$.
A complete simply connected Riemannian manifold of nonpositive sectional curvature is called a Hadamard manifold.
Proposition 4 ([44]).

### v2

@@ Abstract
Our interest lies in developing some efficient methods for minimizing the sum of two geodesically convex functions on Hadamard manifolds, with the aim of improving the convergence of the Douglas-Rachford algorithm in Hadamard manifolds. Specifically, we propose two types of algorithms: inertial and non-inertial algorithms. The convergence analysis of both algorithms is provided under suitable assumptions on algorithmic parameters and the geodesic convexity of the objective functions. This convergence analysis is based on fixed-point theory for nonexpansive operators. We also study the convergence rates of these two methods.
@@ 1 Introduction
The Douglas-Rachford algorithm is among the most celebrated splitting algorithms in convex analysis and operator theory [Combettes2008Inv, Bot2015, Dixit2022, Bergmann2016, ShikherOpt2024, Hieu2018COA, Barshad2023], originally introduced by Douglas and Rachford [Douglas1956] for the numerical solution of partial differential equations. It was later generalized by Lions and Mercier [Lions1967] and independently by Passty [Passty1979] to solve the more general problem of finding a zero of the sum of two maximally monotone operators in Hilbert spaces. Eckstein and Bertsekas [Eckstein1992] provided further insights into its convergence under relaxation schemes and inexact computations.
Beyond operator inclusions, the algorithm also finds wide application in convex optimization, particularly for problems involving the sum of two proper, lower semicontinuous, and convex functions; see [Combettes2008Inv]. Given a Hilbert space $X$ and two proper, convex, lower semicontinuous functions $\phi,\psi:X\to(-\infty,+\infty]$, the Douglas–Rachford algorithm aims to solve the following convex minimization problem:
$\min_{x\in X}\phi(x)+\psi(x).$ (1)
The strength of the Douglas-Rachford algorithm lies in its splitting structure, which allows one to compute the proximal mappings of $\phi$ and $\psi$ separately. This structure makes the algorithm especially useful in large-scale and structured optimization problems [Combettes2008Inv].
Recently, these algorithms were successfully applied in image processing mainly for two reasons: the functional to minimize allows simple proximal mappings within the method and it turned out that the algorithms are highly parallelizable; see [Combettes2008Inv]. Therefore, Douglas-Rachford methods became some of the most popular variational methods for image processing. Continued interest in Douglas-Rachford iteration methods is also due to its excellent performance on various nonconvex problems; see [Bauschke2002, Borwein2011].
Motivated by applications where data naturally reside on nonlinear spaces such as manifolds, common in computer vision, machine learning, and information geometry, there has been growing interest in extending convex optimization methods to manifold settings. Hadamard manifolds, in particular, offer a natural framework for such generalizations due to their geometric properties, including unique geodesics between points [BacakNAA2014]. In the past two decades, several results from linear spaces have been extended to manifold settings because of their applications in many areas of science, engineering, and management (see [Martin2, Martin3, Ferreira2002, Fereira2005]).
@@
$\min_{x\in\mathcal{M}}\phi(x)+\psi(x).$ (2)
To compute the solution to problem (2), Bergmann et al. [Bergmann2016] proposed the following Douglas-Rachford algorithm on Hadamard manifold:
@@
where $\lambda>0$, $\{\alpha_{n}\}$ is a sequence in $[0,1]$ with $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$ and $\gamma(x,y;\cdot)$ is the unique geodesic joining $x$ to $y$ in $\mathcal{M}$. The operators $R_{\lambda\phi}$ and $R_{\lambda\psi}$ denote the reflection of $\phi$ and $\psi$, respectively. Geometrically, $R_{\lambda\phi}(x)$ is obtained by reflecting the point $x$ with respect to its proximal point $\operatorname*{Prox}_{\lambda\phi}(x)$ along the unique geodesic joining them. Bergmann et al. [Bergmann2016] showed that if $R_{\lambda\phi}R_{\lambda\psi}$ is nonexpansive, then the sequence generated by method (3) converges to an element $v$ of $\operatorname*{Fix}(R_{\lambda\phi}R_{\lambda\psi})$ such that $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2).
It is worth noting that by setting $T={R}_{\lambda\phi}{R}_{\lambda\psi}$, the Douglas-Rachford algorithm (3) can be seen as a Riemannian Krasnosel’skiĭ–Mann iteration of the form
@@
which is the natural extension of the classical Krasnosel’skiĭ–Mann scheme from Hilbert spaces to Hadamard manifolds; see [Bergmann2016]. This observation connects optimization and fixed point theory and hence problem (2) can now be viewed as a fixed point problem.
This naturally leads to the framework of fixed point theory. Various problems across diverse domains, such as image reconstruction [Byrne2004], signal processing [Byrne2004], variational inequality problems [Mercier1980], and convex feasibility problems [Bauschke1996] can be formulated as fixed point problems. Let $C$ be a nonempty, closed and convex subset of a real Hilbert space $X$ and $T:C\to C$ be a nonlinear operator. Then, the fixed point problem can be stated as:
@@
To solve such problems, numerous iterative techniques have been developed. One of the most widely used is the Mann iterative method, introduced by Mann [Mann1953] in 1953. The Mann iteration method has its limitations, as it may not effectively approximate the fixed points of pseudocontractive mappings (see [Chidume2001]). To tackle this challenge, Ishikawa [Ishikawa1974] proposed an iterative approach, extensively investigated by multiple authors across different spaces to compute the fixed point of pseudocontractive and nonexpansive mappings (see [Takahashi2008, Acedo2007, Dotson1970]).
Building on this, Agarwal et al. [Agarwal2007] introduced the S-iteration method as follows:
@@
where $\{\alpha_{n}\}$ and $\{\beta_{n}\}$ are sequences in $(0,1)$ with $\sum_{n=1}^{\infty}\alpha_{n}\beta_{n}(1-\beta_{n})=\infty$. In 2011, Sahu [Sahu2011] proposed a modified version of the S-iteration method (5), called the normal S-iteration method, also known as the hybrid Picard-Mann iteration method, which is defined as follows:
@@
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$. Continuing in this direction, Sahu [Sahusoft2020] observed that applying the operator multiple times in each iteration can significantly enhance the convergence. This led to the $p$-accelerated normal S-iteration method [Sahusoft2020], defined as follows:
@@
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ and $p$ is the power of acceleration. For $p=1$, the $p$-accelerated normal S-iteration method reduces to the normal S-iteration method defined by (6). It is observed in [Sahusoft2020] that the $p$-accelerated normal S-iteration method has better convergence performance than the Mann iterative method and normal S-iterative method for the fixed point of nonexpansive operators with convex domain in infinite dimensional Hilbert space. In fact in most cases $p$-accelerated normal S-iteration method works better than the inertial Mann iteration method due to the parameter $p$.
Inertial-type extrapolation techniques play a critical role in accelerating iterative methods. The foundation of such methods can be traced back to Polyak [Polyak1964], who introduced the heavy ball method. Alvarez and Attouch [Alvarez2001] extended the concept to improve the performance of the proximal point algorithm, proposing the inertial proximal point algorithm for solving inclusion problems.
In the existing literature, the utilization of inertial extrapolation-based methods has been thoroughly explored and implemented in various contexts (see [Ochs2014, Bot2016, Chen2014, Samir2023, Attouch2019] and references therein). Recently, researchers have developed several iterative algorithms by incorporating inertial extrapolation techniques, such as the fast shrinkage thresholding algorithm [Beck2009(1)], inertial forward-backward algorithm [Lorenz2015] and inertial Douglas-Rachford splitting algorithm [Bot2015].
Building upon the development of inertial and accelerated iterative methods, Mainge [Mainge2008] proposed a novel approach in 2008 by combining inertial extrapolation with the Mann method, aiming to compute the fixed point of nonexpansive mappings on a real Hilbert space. The inertial Mann algorithm is defined as follows:
@@
where $\{\theta_{n}\}$ and $\{\alpha_{n}\}$ are sequences in $[0,1)$. Mainge [Mainge2008] studied the weak convergence of inertial Mann algorithm (8) under the conditions that $\theta_{n}\in[0,\theta]$ for all $n\in\mathbb{N}$, where $\theta\in[0,1)$, $\sum_{n=1}^{\infty}\theta_{n}\|x_{n}-x_{n-1}\|^{2}<\infty$, and $\inf_{n\in\mathbb{N}}\alpha_{n}>0$ and $\sup_{n\in\mathbb{N}}\alpha_{n}<1$.
The second condition that $\sum_{n=1}^{\infty}\theta_{n}\|x_{n}-x_{n-1}\|^{2}<\infty$, taken by Mainge [Mainge2008] is often difficult to verify in practice. Boţ et al. [Bot2015] proposed a modification in algorithm (8) by replacing this condition with an alternative criterion that the sequence $\{\theta_{n}\}$ is nondecreasing in $0,\theta]$ with $\theta_{1}=0$ and $\theta\in[0,1)$ such that
@@
Recently, Dixit et al. [Dixit2020] proposed the inertial normal S-iteration method for the solution of problem (4) for a nonexpansive operator, which in practice converges faster than the Mann iteration method, inertial Mann iteration method and normal S-iteration method.
The verification of condition (9), as considered by [Bot2015], is a challenging task. To address this issue, Sahu [Sahusoft2020] proposed a new convergence criterion for the inertial Mann iteration. This criterion assumes that the sequences $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ satisfy the following conditions: there exist constants $a,b>0$ such that $0<a\leq\alpha_{n}\leq b<1$ for all $n\in\mathbb{N}$, the sequence $\{\theta_{n}\}$ is nondecreasing in $[0,1)$, and there exists $\theta\in(0,1)$ such that $0\leq\theta_{n}\leq\theta<\frac{\epsilon}{1+\epsilon+\max\{1,\epsilon\}}$ for all $n\in\mathbb{N}$, where $\epsilon=\frac{1-b}{b}$. These conditions are easier to verify in practice.
Motivated by the work of Sahu [Sahusoft2020] and Bergmann et al. [Bergmann2016], the main purpose of this paper is to introduce an inertial Douglas-Rachford method and $p$-accelerated normal S-Douglas-Rachford method for solving problem (2) and study its convergence analysis. Our proposed methods are novel and of interest in the following sense:
  • (i)
    Compared with the classical Douglas–Rachford method on Hadamard manifolds Bergmann et al. [Bergmann2016], we propose two accelerated variants on Hadamard manifolds, namely an inertial Douglas-Rachford method and $p$-accelerated normal S-Douglas-Rachford method. To the best of our knowledge, such inertial and $p$-accelerated Douglas–Rachford schemes have not been previously analyzed in the Riemannian setting.
  • (ii)
    Unlike Sahu [Sahusoft2020] and other works on $p$-accelerated normal S-iteration in Hilbert spaces, we extend the $p$-acceleration idea to Douglas–Rachford splitting methods on Hadamard manifolds. Although some of the conditions are similar to the Hilbert space case, this extension is not straightforward because Hadamard manifolds do not have a linear structure. It requires careful use of geodesic convexity, exponential maps, and comparison geometry.
  • (iii)
    Compared with the inertial Douglas–Rachford and inertial Mann schemes studied in Hilbert spaces by Boţ et al. [Bot2015] and Mainge [Mainge2008], our analysis is carried out entirely in the nonlinear setting of Hadamard manifolds. Moreover, we use different parameter conditions, which are easier to verify in practice than those in Boţ et al. [Bot2015] and Mainge [Mainge2008].
  • (iv)
    We analyze the convergence rates of our proposed algorithms. Specifically, the $p$-accelerated normal S-iteration method converges at a rate of order $o(1/\sqrt{n})$, while the inertial Mann iteration method converges in the minimal residual sense, also at a rate of order $o(1/\sqrt{n})$.
We give an application of these algorithms to minimizing functionals containing multiple summands by introducing a parallel inertial Douglas-Rachford method and parallel $p$-accelerated Douglas-Rachford method. We also apply our parallel Douglas-Rachford methods to solve the generalized Heron problems in Hadamard manifolds and provide numerical experiments.
@@ 2 Preliminaries
Let $C$ be a nonempty subset of a metric space $(X,d)$ and let $T\colon C\to C$ be an operator. Then $T$ is said to be nonexpansive if $d(Tx,Ty)\leq d(x,y)\text{ for all }x,y\in C.$ We denote the set of fixed points of an operator $T\colon C\to C$ by $\operatorname*{Fix}(T)$, thus, $\operatorname*{Fix}(T)=\{x\in C\colon Tx=x\}.$
@@
$R_{T,\{x_{n}\}}(n)=\min_{1\leq i\leq n}d(x_{i},Tx_{i})\text{ for all }n\in\mathbb{N}.$
Observe that $R_{T,\{x_{n}\}}(n)\leq d(x_{i},Tx_{i})$ for all $i=1,2,\ldots,n$ and $\{R_{T,\{x_{n}\}}(n)\}$ is decreasing. We use the notation $o$ to mean that $s_{n}=o(1/t_{n})$ if and only if $\lim_{n\to\infty}s_{n}t_{n}=0.$
Lemma 1 ([Dong2015comments]).
Let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences of positive real numbers such that $\sum_{n=1}^{\infty}a_{n}b_{n}<\infty.$ Suppose that $\sum_{n=1}^{\infty}a_{n}=\infty$ and that the sequence $\{b_{n}\}$ is decreasing. Then $b_{n}=o\!\left(\frac{1}{\sum_{i=1}^{n}a_{i}}\right).$
Definition 2 ([SahuNFAO]).
Let $C$ be a nonempty subset of a metric space $(X,d)$ and let $T\colon C\to C$ be an operator with $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence in $C$. We say that $\{x_{n}\}$ has
@@
    the LEAF point property if $\{x_{n}\}$ satisfies both LE property and AF property.
Lemma 3 ([Sahusoft2020]).
@@
$a_{n+1}\leq a_{n}+\theta_{n}(a_{n}-a_{n-1})+b_{n}\text{ for all }n\in\mathbb{N},$
$\sum_{n=1}^{\infty}b_{n}<\infty$ and there exists a real number $\theta$ with $0\leq\theta_{n}\leq\theta<1$ for all $n\in\mathbb{N}$. Then, the following hold:
@@
For the basic definitions and results regarding the geometry of Riemannian manifolds, we refer to [Docarmo, Sakai, Udriste].
Let $\mathcal{M}$ be an $m$-dimensional differentiable manifold and let $x\in\mathcal{M}$. The set of all tangent vectors at the point $x$, denoted by $\mathcal{T}_{x}\mathcal{M}$, is called the tangent space of $\mathcal{M}$ at $x$. The tangent bundle of $\mathcal{M}$ is defined as the disjoint union $\mathcal{T}\mathcal{M}=\bigcup_{x\in\mathcal{M}}\mathcal{T}_{x}\mathcal{M}$. A differentiable manifold $\mathcal{M}$ with a Riemannian metric $\langle\cdot,\cdot\rangle$ is said to be a Riemannian manifold. The induced norm on $\mathcal{T}_{x}\mathcal{M}$ is denoted by $\|\cdot\|_{x}$. The Riemannian distance between $x$ and $y$ in $\mathcal{M}$ is designated by $d(x,y)$. A geodesic joining $x$ to $y$ in the Riemannian manifold $\mathcal{M}$ is said to be a minimal geodesic if its length is equal to $d(x,y)$. We denote geodesic joining $x$ and $y$ by $\gamma(x,y;\cdot)$, i.e., $\gamma(x,y;\cdot)\colon[a,b]\to\mathcal{M}$ is such that, for $a,b\in\mathbb{R}$, we have $\gamma(x,y;a)=x$ and $\gamma(x,y;b)=y$.
@@
Let $\nabla$ be the Riemannian connection associated with the Riemannian manifold $\mathcal{M}$. The parallel transport on the tangent bundle $\mathcal{T}\mathcal{M}$ along the geodesic $\gamma(x,y;\cdot)$ with respect to Riemannian connection $\nabla$ is defined as $P_{\gamma(x,y;b),\gamma(x,y;a)}\colon\mathcal{T}_{\gamma(x,y;a)}\mathcal{M}\to\mathcal{T}_{\gamma(x,y;b)}\mathcal{M}$ such that $P_{\gamma(x,y;b),\gamma(x,y;a)}(v)=V(\gamma(x,y;b))\text{ for all }a,b\in\mathbb{R},v\in\mathcal{T}_{x}\mathcal{M},$ where $V$ is the unique vector field satisfying $\nabla_{\gamma^{\prime}(x,y;t)}V=0$ for all $t\in[a,b]$ and $V(\gamma(x,y;a))=v$. A Riemannian manifold $\mathcal{M}$ is said to be complete if for any $x\in\mathcal{M}$ all geodesic emanating from $x$ are defined for all $t\in\mathbb{R}$. By Hopf-Rinow Theorem [Sakai], if $\mathcal{M}$ is complete, then any pair of points in $\mathcal{M}$ can be joined by a minimal geodesic. Moreover, $(\mathcal{M},d)$ is a complete metric space. A complete simply connected Riemannian manifold of nonpositive sectional curvature is called a Hadamard manifold. Since in a Hadamard manifold $\mathcal{M}$, any two points can be joined by a unique geodesic, we denote the parallel transport by $P_{y,x}$ instead of $P_{\gamma(x,y;b),\gamma(x,y;a)}$. If $\mathcal{M}$ is a Hadamard manifold, then the exponential map $\exp_{x}\colon\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ at $x\in\mathcal{M}$ is defined by $\exp_{x}v=\gamma_{x,v}(1)$ for all $v\in\mathcal{T}_{x}\mathcal{M}$, where $\gamma_{x,v}\colon\mathbb{R}\to\mathcal{M}$ is a unique geodesic starting from $x$ with velocity $v$, i.e., $\gamma_{x,v}(0)=x$ and $\gamma^{\prime}_{x,v}(0)=v$. It is known that $\exp_{x}(tv)=\gamma_{x,v}(t)$ for each real number $t$. For the exponential map, the following properties hold:
@@
The following proposition establishes the relationship between geodesics and the exponential map.
Proposition 5 ([Sakai]).